For thousands of years, gold has been considered one of the Earth's most precious metals. One hundred percent pure gold is 24 karat gold, which is too soft to be made into jewelry. Most gold jewelry is 14-karat gold, approximately 58% gold. If 18-karat gold is 75% gold and 12-karat gold is 50% gold, how much of each should be used to make a 14-karat gold bracelet weighing 300 grams?

Answer by mananth(16946)

percent ---------------- quantity
Gold type I 75.00% ---------------- x gm
Gold type II 50.00% ------ 300 - x gm
Mixture 58.00% ---------------- 300
300
75.00% x + 50.00% ( 300 - x ) = 58.00% * 300
75 x + 50 ( 300 - x ) = 17400
75 x + 15000 - 50 x = 17400
75 x - 50 x = 17400 - -15000
25 x = 2400
/ 25
x = 96 gm 75.00% Gold type I ( 18 Karat)
204 gm 50.00% Gold type II ( 12 karat)

m.ananth@hotmail.ca